### v1

# A posteriori closure of turbulence models: are symmetries preserved ?

Abstract
Turbulence modeling remains a longstanding challenge in fluid dynamics. Recent advances in data-driven methods have led to a surge of novel approaches aimed at addressing this problem. This work builds upon [1], where we introduced a new closure for a shell model of turbulence using an a posteriori (or solver-in-the-loop) approach. Unlike most deep learning-based models, our method explicitly incorporates physical equations into the neural network framework, ensuring that the closure remains constrained by the underlying physics benefiting from enhanced stability and generalizability. In this paper, we further analyze the learned closure, probing its capabilities and limitations. In particular, we look at joint probability density functions to assess whether cross-correlations are well preserved or if just the mean behavior is captured. Additionally, we investigate the scale invariance of multipliers—ratios between adjacent shells—within the inertial range. Although our model excels in reproducing high-order observables such as flatness, it breaks this known symmetry near the cutoff, indicating a fundamental limitation. We discuss the implications of these findings for subgrid-scale modeling in 3D turbulence and outline directions for future research.
1 Introduction
Turbulence is ubiquitous in nature, from the coffee we stir in the morning to the primordial states of our universe [2]. Despite its ubiquity, understanding and modeling turbulence remains one of the most challenging tasks in classical physics. While we have known its governing equations for over a century, the full phenomenology of turbulence remains elusive. Simulating turbulence numerically is also very computationally heavy. The degrees of freedom scale as a power law of the Reynolds number, $Re=uL/\nu$, which means that studying very high Reynolds number turbulence remains impossible for the most part. One approach to reduce the degrees of freedom is Large Eddy Simulation (LES) [3], where a filter is applied, typically within the inertial range, to resolve only the largest, most energy-containing eddies. However, to close the LES system, the effects of unresolved subgrid scales must be modeled. This constitutes the LES modeling problem. This problem is often addressed phenomenologically by introducing an eddy diffusivity term to stabilize the system. These solutions, albeit useful for practical application, fail to capture the correct subgrid scale statistics. Moreover, although decades of research have been spent in LES modeling, there are still very fundamental open questions in the field. Can a subgrid-scale closure be formulated that accurately captures the statistics of the resolved scales, especially in high Reynolds number flows where intermittency plays a crucial role? In turbulence, velocity fields are not statistically self-similar due to intermittency, making it essential for closures to account for the multifractal nature of the subgrid scales.
Recently, deep learning (DL) approaches have gained attention for addressing the LES modeling problem. DL excels at approximating functions with unknown functional forms. By learning from data, DL methods can, albeit in a black-box manner, accurately approximate high-dimensional functions.
To investigate the applicability of deep learning in this context, we use shell models of turbulence [4]. Shell models offer a simplified representation of 3D homogeneous isotropic turbulence (HIT), providing a tractable alternative to the full Navier-Stokes equations. Despite their simplicity, shell models preserve key turbulence phenomena, such as energy cascade, intermittency, anomalous scaling, and multifractal energy dissipation.
Our primary focus is on the ability of subgrid-scale (SGS) models to accurately capture extreme events and rare fluctuations, which demand high precision, extensive statistical sampling, and long inertial ranges. Learning SGS models from data requires long temporal evolutions, presenting a fundamental challenge for LES of three-dimensional turbulence. This challenge is currently best addressed through shell models. Although LES of three-dimensional HIT remains the ideal test case, existing machine learning-based LES tools are hindered by their low resolution and the proximity of the cutoff to the forcing scales, preventing the emergence of anomalous behavior. Shell models, in contrast, provide a more suitable framework for rigorously testing the effectiveness of LES models under conditions of strong intermittency and for assessing whether solver-in-the-loop approaches can improve the tracking of extreme fluctuations over time. Furthermore, higher-order statistics—particularly of fourth order and beyond—naturally involve non-local interactions in Fourier space, implying that SGS fluctuations can influence the resolved scales in non-trivial ways. This observation is critical when aiming for accuracy beyond second-order moments, which are typically the focus of LES validation. Shell models thus offer a controlled environment to probe whether a closure model can correctly capture such high-order interactions and preserve the cross-scale dynamics.
Several previous studies have explored closures within the shell model framework. Biferale et al. [5] initially approached the closure problem phenomenologically, using Kolmogorov multipliers. More recently, Ortali et al. [6] employed deep learning to learn the closure, while Domingues Lemos [7, 8] integrated probabilistic machine learning tools within the hidden symmetric framework of shell models [9] to address the closure. Building on these developments, we have also contributed to the field [1] through a solver-in-the-loop approach [10] for shell model closure based on the differentiable physics paradigm. This method unrolls the system in time during training, allowing the neural network to interact with a numerical solver for the governing equations over an arbitrary number of time steps before the loss function is computed. Rather than attempting to match the instantaneous, a priori evolution of the ground truth, we adopt an a posteriori approach, where the network learns how its errors propagate through time while receiving more realistic inputs. The induced closure was analyzed by examining high-order moments such as flatness up to the tenth order, demonstrating good agreement with the ground truth and unconditionally stable behavior.
This paper further extends this investigation, looking deeper into the learned closure. We will mainly focus on correlations between resolved and unresolved (i.e., predicted by the neural networks) variables, to assess whether the model accurately captures the cross-correlations, or whether it is only able to capture the mean behavior. Moreover, we will look into Kolmogorov multipliers statistics. These are ratios of velocity fluctuations between adjacent shells and have been shown to have universal statistics in the inertial range of fully developed turbulence. They were introduced by Kolmogorov, who hypothesized their self-similar statistics and scale invariance [11]. We will look at whether the scale invariance of these multipliers is broken by the subgrid-scale model, or whether the learned closure is able to preserve this symmetry, despite being trained purely on an $L^{2}$ norm of the velocity fields.
The remainder of the paper is structured as follows. In Section 2, we introduce the shell model framework and the mathematical formulation of our problem. Section 3 details the solver-in-the-loop approach and the methodology for learning an a posteriori closure. In Section 4, we present our results, focusing on statistical properties and preservation of known symmetries. Finally, we discuss our findings in Section 5 and outline potential directions for future research.
2 Shell models
Shell models are a class of turbulence models. They model homogeneous isotropic turbulence in Fourier space, considering only sparse coupling between the different shells. We discretize the 1D Fourier space with a geometric progression of shells that are equispaced on a logarithmic lattice, $k_{n}=k_{0}\lambda_{n}$, where one typically uses $\lambda\!=\!2$ and $k_{0}\!=\!1$. We consider the Sabra shell model [12], where the set of ordinary differential equations governing the dynamics of the different shells is given by:
$\frac{du_{n}}{dt}=i\Big{(}ak_{n+1}u_{n+2}u_{n+1}^{*}+bk_{n}u_{n+1}u_{n-1}^{*}-ck_{n-1}u_{n-1}u_{n-2}\Big{)}-\nu k_{n}^{2}u_{n}+f_{n}\,,$ (1)
where $u_{n}\!\in\!\mathbb{C}$, represents the velocity fluctuations at a given shell $n$. The free parameters $a$, $b$, and $c$ are chosen such that $a+b+c=0$, which corresponds to conservation of energy $E\!=\!\sum_{n}|u_{n}|^{2}$ and helicity $H\!=\!\sum_{n}(a/c)^{n}|u_{n}|^{2}$ in the inviscid, unforced limit, mimicking the quadratic inviscid invariants in the Euler equations.
The first term in the RHS of Equation 1 represents the non-linear term, where we can see that only nearest neighbour and next-to-nearest neighbour interactions are preserved, making this model local in Fourier space; the second term represents the viscous term; and the third term represents large-scale forcing.
The left boundary conditions are $u_{-1}\!=\!u_{-1}\!=\!0$. For the fully resolved system, we integrate Equation 1 for shells $n\!=\!0,\ldots,N$, where $N$ is chosen with some margin of the dissipative cutoff scale, $N_{\eta}$, as to fully resolve all of the dynamics. In the case of closing the system in an LES fashion, we have to consider a cutoff shell $N_{c}$, after which we only have to model the shells $N_{c}\!+\!1$ and $N_{c}\!+\!2$ (the left BCs remain the same), as depicted in Figure 1. The velocity fields of these shells will appear in the equations for $u_{N_{c}-1}$ and $u_{N_{c}}$:
$\frac{du_{N_{c}}}{dt}$ $=i\Big{(}k_{N_{c}+1}u_{N_{c}+2}u_{N_{c}+1}^{*}-\frac{1}{2}k_{N_{c}}u_{N_{c}+1}u_{N_{c}-1}^{*}+\frac{1}{2}k_{N_{c}-1}u_{N_{c}-1}u_{N_{c}-2}\Big{)}-\nu k_{N_{c}}^{2}u_{N_{c}}$ (2a)
$\frac{du_{N_{c}-1}}{dt}$ $=i\Big{(}k_{N_{c}}u_{N_{c}+1}u_{N_{c}}^{*}-\frac{1}{2}k_{N_{c}-1}u_{N_{c}}u_{N_{c}-2}^{*}+\frac{1}{2}k_{N_{c}-2}u_{N_{c}-2}u_{N_{c}-3}\Big{)}-\nu k_{N_{c}-1}^{2}u_{N_{c}-1}$ (2b)
Figure 1: Energy spectrum $|u_{n}|^{2}$ as a function of the wavenumber $k_{n}$. The red-circled markers indicate the reduced system, while black crosses represent the fully resolved system. The inset zooms into the cutoff region, highlighting the nonlinear energy transfer through triadic interactions, with blue and green arrows indicating interactions for the two shells preceding the cutoff. It becomes clear that to close the system, only the two shells after the cutoff are needed to model. The x-axis is labeled in terms of $\log_{2}k_{n}$, with markers for the Kolmogorov scale $k_{\eta}$ and the maximum resolved wavenumber $k_{N}$.
3 A posteriori closure
To close the system, we use a deep neural network that predicts the two missing shells, $u_{N_{c}+1}$ and $u_{N_{c}+2}$, using as input the three preceding shells. Unlike an a priori approach, which trains the model to minimize one-step errors, we follow an a posteriori or solver-in-the-loop strategy, where the neural network operates within the evolving simulation. This allows it to learn not only accurate short-term predictions but also how its own errors propagate over time and influence the system’s long-term behavior.
A common challenge in machine learning is the distribution shift, i.e., when a model, after making several predictions, encounters states it has never been seen during training, leading to increasing errors. This is well known in autoregressive generative models, such as language models predicting text sequences [13]. Similarly, in physics-based modeling, an imperfect closure leads to accumulated deviations that push the system into untrained regions of phase space. By embedding the neural network directly within the governing equations, we ensure it learns to operate under realistic conditions, mitigating the effects of distribution shift. This approach leads to improved stability, better generalization, and a more physically consistent closure.
Let the reduced system $(\tilde{u}_{0},\tilde{u}_{1},\dots,\tilde{u}_{N_{c}})$ reside in the LES manifold $\mathcal{S}$, while the full system $(u_{0},u_{1},\ldots,u_{N})$ spans the reference manifold $\mathcal{R}$. The governing equations of the shell model define a dynamical system on $\mathcal{R}$, but large eddy simulations truncate the system to $\mathcal{S}$, introducing closure errors. To bridge this gap, we introduce a correction operator $\mathcal{C}:\mathcal{S}\!\rightarrow\!\bar{\mathcal{R}}$ parameterized by a neural network, where $\bar{\mathcal{R}}$ represents the truncated reference manifold, restricted to the resolved scales $(u_{0},u_{1},\ldots,u_{N_{c}})$. This operator predicts the unresolved shells $\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}=\mathcal{C}(\tilde{u}_{N_{c}-2},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}})$, enabling reconstruction of the full state in $\mathcal{C}$, via an explicit time integration of the missing terms: this corresponds to an additive correction to the resolved terms that are integrated through a fourth-order Runge-Kutta scheme.
The operator $\mathcal{C}$ is implemented as a fully connected feedforward neural network (multi-layer perceptron, MLP) with $L$ hidden layers. Given an input vector
$\bm{x}=(\tilde{u}_{N_{c}-2},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}})\in\mathbb{C}^{3},$
the output $\tilde{\bm{u}}=(\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2})\in\mathbb{C}^{2}$ is computed through a sequence of affine transformations and nonlinear activations:
$\bm{h}^{(0)}=\bm{x},$
$\bm{h}^{(\ell)}=\sigma\left(\bm{W}^{(\ell)}\bm{h}^{(\ell-1)}+\bm{b}^{(\ell)}\right),\quad\text{for }\ell=1,\dots,L,$
$\tilde{\bm{u}}=\bm{W}^{(L+1)}\bm{h}^{(L)}+\bm{b}^{(L+1)}.$
Here, $\bm{W}^{(\ell)}$ and $\bm{b}^{(\ell)}$ are the weight matrix and bias vector for the $\ell$-th layer, while $\sigma(\cdot)$ is a nonlinear activation function, in our case ReLU. The final layer has a linear activation to ensure unrestricted output values.
The evolution of the system is governed by a discrete solver $\mathcal{P}$, which advances the resolved shells via:
$\tilde{\bm{u}}(t+\Delta t)=\mathcal{P}(\tilde{\bm{u}}(t),\tilde{u}_{N_{c}+1},\tilde{u}_{N_{c}+2}),$
where $\tilde{\bm{u}}(t)=(\tilde{u}_{0},\dots,\tilde{u}_{N_{c}})_{t}$. Unlike a priori closure models trained to minimize single-step errors, our a posteriori approach integrates $\mathcal{C}$ recurrently into $\mathcal{P}$.
During training, the solver interacts with $\mathcal{C}$ over $m$ steps, exposing it to its own error dynamics. The loss function evaluates deviations from the reference trajectory across this horizon:
$\mathcal{L}(\tilde{\bm{u}}^{\bm{\theta}},\bm{u}^{\text{ref}})=\sum_{k=1}^{m}\left\|\tilde{\bm{u}}^{\bm{\theta}}(t+k\Delta t)-\bm{u}^{\text{ref}}(t+k\Delta t)\right\|_{2}^{2},$ (3)
where $\bm{u}^{\text{ref}}$ is obtained from a direct numerical simulation of $\mathcal{R}$. The optimization problem is then formulated as:
$\min_{\bm{\theta}}\mathcal{L}(\tilde{\bm{u}}^{\bm{\theta}},\bm{u}^{\text{ref}}),$
where $\bm{\theta}=\{\bm{W}^{(\ell)},\bm{b}^{(\ell)}\}_{\ell=1}^{L+1}$ represents the neural network parameters. This multi-step optimization ensures $\mathcal{C}$ learns to mitigate error accumulation and distribution shift, as illustrated in Figure 2.
Figure 2: Sketch illustrating the solver-in-the-loop approach. The LES trajectory $\tilde{\bm{u}}$ is evolved over $m$ time steps using the discrete solver $\mathcal{P}$, while the correction operator $\mathcal{C}$, parameterized by a neural network, predicts the unresolved scales. The objective is to iteratively refine $\mathcal{C}$ during training, bringing $\tilde{\bm{u}}$ closer to the reference trajectory $\bm{u}^{\text{ref}}$, which integrates the full system dynamics from the same initial condition. Here, the subscript denotes the time step index of the trajectory, i.e., $\bm{u}_{k}=\bm{u}(t+k\Delta{t})$, and not the shell index.
4 Results
In this section, our DL-closed reduced system of equations, hereon referred to as LES-NN, is evaluated by comparing its predictions against the fully resolved ground truth across multiple statistical measures. We analyze its performance in reproducing key turbulence characteristics, including structure functions, temporal evolution of energy, and scale-dependent correlations. Particular attention is given to the ability of the model to capture the tails of joint probability density functions between resolved and unresolved variables; as well as scale invariant statistics following the third hypothesis of K62 theory, and whether this symmetry is preserved by the closure.
The structure functions, defined as the difference between velocity increments across different scales to some power $p$, can be defined for the shell models as
$S_{n}^{(p)}=\langle|u_{n}|^{p}\rangle\,.$ (4)
These are shown in Figure 3 up to the cutoff shell, for $p=1,\ldots,6$ for both the fully resolved system (ground truth) and the LES-NN model. There is a great agreement between the model and the ground truth for all order structure functions showed, even close to the cutoff scale, which is the subgrid scales effect is more noticeable.
Figure 3: Structure functions $S_{n}^{(p)}$ as a function of shell index $n$ for different orders $p=1,\ldots,6$. The LES-NN model is compared with the ground truth solution.
In Figure 4, we show the time evolution of the logarithm of the normalised velocity field (velocity field scales as $u_{n}\!\sim\!\smash{k_{n}^{-1/3}}$) for both the ground truth (GT) and the LES-NN model, along with the normalized error between the two. Both the GT and LES-NN models start from the same initial condition (IC). It is evident that the LES-NN model rapidly decorrelates from the GT. This behavior is expected, as we cannot synchronize past the Lyapunov time. For this particular IC, the discrepancy is particularly clear because we have chosen an IC that is within a burst of energy that has already propagated to the small scales, where the dynamics are much faster. As a result, the effective Lyapunov time aligns with these faster time scales.
Figure 4: Time evolution of the normalised energy density $\log_{2}(|u_{n}|^{2}k_{n}^{2/3})$ for (a) the ground truth and (b) the LES-NN model. (c) Absolute normalised error $\left||u_{n}|^{2}-|\tilde{u}_{n}|^{2}\right|k_{n}^{2/3}$ between the two fields. The x-axis represents the wavenumber index $n$, while the y-axis represents the normalized time $t/\tau_{0}$. The dashed line represents the cutoff scale.
Figure 5 shows the joint PDF between the sum of the resolved energy of the shells used as input for the neural network, ${\tilde{u}_{N_{c}},\tilde{u}_{N_{c}-1},\tilde{u}_{N_{c}-2}}$, and the sum of the unresolved energy, i.e., the sum of the energies of the shells that the NN outputs, ${\tilde{u}^{\theta}_{N_{c}+1},\tilde{u}_{N_{c}+2}^{\theta}}$, for both the ground truth and the LES-NN model. While the overall shape of the PDF is correctly predicted by the model, it appears to primarily capture the mean behavior rather than accurately reproducing the full distribution.
There are a few potential reasons for this discrepancy. One possible cause is the architecture of the neural network itself (multi-layer perceptron), which may lack the necessary complexity to fully capture the higher-order correlations in the data. Spectral bias, a tendency of neural networks to favor learning smooth, low-frequency components over more complex high-frequency variations, could further contribute to the underestimation of the tails in the predicted distribution. Additionally, it can also be an outcome of the choice of the loss function as an MSE between velocity fields of the resolved shells, thus not including the unresolved ones explicitly.
Figure 5: Contours of the joint PDF of the summed energy across the input shells for the neural network (x-axis) and the subgrid-scale output shells (y-axis) for both the ground truth and the LES-NN model.
In his seminal 1962 study [11], Kolmogorov advanced a theoretical framework to characterize the intermittency observed in turbulence at small scales. Central to this framework was the concept of multipliers, defined as scale-dependent ratios of velocity increments. These multipliers, denoted as $w_{ij}(x;\ell,\ell^{\prime})$, were expressed as:
$w_{ij}(x;\ell,\ell^{\prime})=\frac{\delta_{i}v_{j}(x,\ell)}{\delta_{i}v_{j}(x,\ell^{\prime})}\,,$ (5)
where $\delta_{i}v_{j}(x,\ell)=v_{j}(x+\ell\mathbf{e}_{i})-v_{j}(x)$ represents the increment of the $j$-th velocity component across a spatial separation $\ell$ along the unit vector $\mathbf{e}_{i}$. Kolmogorov posited that at sufficiently high Reynolds numbers, the statistical distributions of these multipliers would exhibit universality—dependent solely on the scale ratio $\ell/\ell^{\prime}$ and independent of the absolute scales $\ell$ or $\ell^{\prime}$. Furthermore, he hypothesized statistical independence between multipliers associated with scales separated by large intervals, a conjecture later termed his third hypothesis [14].
For shell models, one can easily define amplitude and phase multipliers [15], respectively, via:
$w_{n}(t)=\left|\frac{u_{n}(t)}{u_{n-1}(t)}\right|\,,$ (6)
$\Delta_{n}(t)=\theta_{n}(t)-\theta_{n-1}(t)-\theta_{n-2}(t)\,,$ (7)
where $\theta_{n}=\arg u_{n}$.
In Figure 6, we show the PDF of the amplitude and phase multipliers for the ground truth and our model, LES-NN, for shells $n\!=\!6,\ldots,N_{c}$. For the fully resolved model, we see the expected collapse. However, for the LES-NN model, the learned closure breaks this symmetry, and the multipliers are no longer scale invariant near the cutoff. This is a limitation of the model.
Figure 6: PDF of the (a) amplitude multipliers, $w_{n}$, and (b) phase multipliers, $\Delta_{n}$, for the ground truth (GT) and LES-NN model, for the shells $7$ to $N_{c}$.
We extend this analysis in Figure 7 by examining the joint PDF of the amplitude of the product of the last two resolved multipliers and the product of the two unresolved multipliers (which include the neural network’s predictions). The LES-NN model accurately captures the resolved multipliers but fails significantly in the unresolved ones, leading to incorrect predictions not only in the tails but also in the mean. As in all highly nonlinear, high-dimensional optimization problems, the neural network training converged to a local minimum—sufficient to reproduce high-order structure functions correctly but learning a flawed way to match the subgrid-scale statistics, resulting in good overall results but with clear failures in certain observables. This suggests room for improvement, possibly in the choice of loss function, input-output representation, or overall problem formulation. We leave this as an open question for future investigation.
Figure 7: Contours of the joint PDF between the product of the last two resolved multipliers (x-axis) and the product of the two subgrid scale multipliers (y-axis) for both the ground truth and the LES-NN model.
Following Eyink et al. [16], we can further define:
$\sigma_{n}(t)=\ln w_{n}(t)\,,$ (8)
$U_{n}(t)=\exp(i\Delta_{n}(t))\,,$ (9)
where $\sigma_{n}(t)$ is a “local slope” of the multiplier $w_{n}(t)$ and $U_{n}(t)$ is a 2D rotator spin that lives on the unit circle in the complex plane. We are interested in studying the correlation of these variables across different shells. For this purpose, we define the following correlation function [16]:
$C_{XY}(n,m)=\left\langle X_{n}^{*}Y_{m}\right\rangle-\left\langle X_{n}^{*}\right\rangle\left\langle Y_{m}\right\rangle\,.$ (10)
The results for $X\!=\!Y\!=\!\sigma$ and $X\!=\!Y\!=\!U$ shown in Figure 8 for the GT and LES-NN model. Shell $n\!=\!11$ is kept as the reference to compute correlations. We see the expected fast decay with increasing $|m-n|$; however, the correlations near the cutoff are wrongly estimated for the LES-NN model. This helps shine a light on the lack of collapse of the multipliers near the cutoff.
Figure 8: Spin-spin correlation functions with shell $n\!=\!11$ as reference. The dashed line represents the cutoff scale.
5 Discussion
In this work, we have delved deeper into the performance of a closure trained using a solver-in-the-loop approach for closure of the shell models of turbulence [1]. We focused on correlations between resolved and unresolved variables, as well as universal single-time statistics of the inertial range of fully developed turbulence, the so-called Kolmogorov multipliers (ratio between adjacent shells). We showed that, even though our closure is able to reproduce high-order moments, such has structure functions of $p=6$ order with great accuracy, it still underpredicts the cross-correlations between the energy of the resolved and unresolved shells. While the overall shape of the joint PDF is well captured, the learned closure tends to focus on the mean behavior. This could be attributed to several factors. One potential reason is the architecture of the neural network in use (i.e., a simple multi-layer perceptron), which might lack the complexity needed to fully capture higher-order correlations. Additionally, the spectral bias inherent in neural networks could cause them to prioritize smoother, low-frequency components, leading to an underestimation of the tails in the predicted distribution. The choice of the loss function, which focuses solely on the MSE between velocity fields of the resolved shells, may also contribute, as it does not explicitly account for the unresolved shells.
Regarding the performance of the learned closure in terms of the multipliers, the LES-NN model breaks their scale invariance near the cutoff: the PDFs of the multipliers near the cutoff no longer collapse. The perfect LES model would preserve this symmetry in the inertial range. Our model was not trained using multipliers, so it is not completely surprising. The interesting question is whether training the model focusing on multiplier statistics in the loss function will yield a better-performing model than one trained on simple velocity differences. This was explored by Domingues Lemos et al. in [7] and more recently in [8] by working in the hidden symmetry framework [9]. Compared to these, our closure performs better on observables such as structure functions and flatness, but not in terms of multipliers statistics. Training with a solver-in-the-loop approach and within the multipliers framework might result in a better-learned closure. Another powerful benefit of using the multipliers closure is that, in principle, the learned closure should be cutoff independent (given that the cutoff is not affected by the forcing or the dissipation scales). This was numerically demonstrated by Domingues Lemos in [8].
We report these results in the hope of advancing our understanding of turbulence closure problems, particularly in relation to the optimal setup, which has strong implications for both data-driven based and phenomenology based closures. Additionally, the study on multiplier statistics, and the evidence that our model breaks this symmetry are potential signs that the current framework is not ideal, and one that explicitly contains or enforces this scale invariance might be more attractive. This has potential implication in subgrid scale modeling in 3D Navier-Stokes turbulence, as a similar rescaling procedure can be done, leading to scale invariance [17]. Naturally, additional challenges are introduced, but the overall message might still remain true. Future studies on the topic are, therefore, needed.
Acknowledgments
The authors benefited from discussions with G. Eyink, A. Mailybaev, M. Sbragaglia. This research was supported by European Union’s HORIZON MSCA Doctoral Networks programme under Grant Agreement No. 101072344, project AQTIVATE (Advanced computing, QuanTum algorIthms and data-driVen Approaches for science, Technology and Engineering), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme Smart-TURB (Grant Agreement No. 882340), and through an Inria Chair.